# The AI Memory Gap: Users Misremember What They Created With AI or Without

### v1

@@ 3.4.1. Phase 1 – Creating Ideas and Elaborations
The task description displayed the problem text and instructed participants to create five ideas each consisting of 1-3 keywords. These ideas were entered into five input fields displayed on the same page (Figure 3 A).
In the withAI condition, a chatbot interface was available on the right side of the screen (Figure 3 C). Participants could copy text from the chatbot and paste it into the input fields using the usual keyboard shortcuts or buttons located next to each field. In the noAI condition, the chatbot was locked (Figure 3 D) and pasting into the input fields was disabled.
@@
Following this, participants were asked to elaborate on each of their previously submitted ideas with a one-sentence elaboration. Each such elaboration task displayed the task instruction, the original problem statement, and one of the participant’s earlier ideas alongside a text input field (Figure 3 B).
@@
For both ideation and elaboration, the overall progress was shown at bottom of the application (Figure 3 E). After completing elaborations for all 40 ideas, the study concluded with a thank-you screen. Participants were reminded that they would receive an invitation to Phase 2 of the study six days later and that this follow-up phase would then need to be completed within 48 hours.
@@ 3.4.2. Phase 2 - Memory and Source Attribution
After the six-day delay, participants who completed Phase 1 received an invitation to take part in Phase 2. They were presented with the study information and were again required to provide informed consent before proceeding.
We first asked participants to guess the purpose of the study in a free-text entry field. The main part of Phase 2 then consisted of the memory and source attribution tasks (Figure 4). Besides the overall task description, each task page displayed the original problem statement and a solution (elaboration) text created by the participant in Phase 1. The underlying ideas were not shown.
Participants first indicated whether they remembered working on the displayed solution at all. If they did, they were asked: (1) Whether they came up with the underlying idea on their own or with AI support, (2) How confident they were in that attribution, (3) Whether they wrote the solution text on their own or with AI support, and (4) How confident they were in that second attribution. If they did not remember working on the displayed solution at all, they skipped the attribution questions and proceeded to the next memory task.
@@ 3.5. Data sharing
We make our experimental software, study material, analysis code, and the collected data accessible through our project repository: https://osf.io/anonymous/. Two researchers went through all the data collected via free text input fields to ensure that all potentially privacy-sensitive information were removed before publication.
@@ 4.1. Source Attribution Accuracy
Figure 5 shows the model-based interaction plots of predicted source attribution accuracy for all three memory types across conditions (noAI vs. withAI).
@@ 4.1.1. Remembering the Item
Pairwise contrasts (Holm-adjusted) showed that the no-AI workflow (noAI/noAI) yielded higher accuracy than any condition involving AI: vs. withAI/noAI, OR =2.02=2.02; vs. noAI/withAI, OR =3.29=3.29; vs. withAI/withAI, OR =8.15=8.15, all $p<.001$. Among the AI-involved conditions, withAI/noAI exceeded noAI/withAI (OR =1.99=1.99) and withAI/withAI (OR =4.01=4.01); noAI/withAI also exceeded withAI/withAI (OR =2.45=2.45); again all $p<.001$. These results indicate that both AI idea generation and AI elaboration independently impair memory, and their joint use compounds this effect, with the strongest decrement observed when AI contributed at the elaboration stage.
@@ 4.1.2. Remembering the Source of Ideas
Both main effects were negative and significant, qualified by a strong positive interaction: $\beta_{\text{idea({withAI}{})}}=-3.05$, $\beta_{\text{elab({withAI}{})}}=-1.97$, $\beta_{\text{idea({withAI}{})}\times\text{elab({withAI}{})}}=3.81$; all $p<.001$. That is, using AI for the idea alone sharply reduced correct attribution (OR =0.05=0.05), as did using AI only for elaboration (OR =0.14=0.14); when both idea and elaboration came from AI, the interaction increased the odds ($\times 45.3$), partially offsetting the main-effect penalties.
Model-based predicted probabilities (Table 2) make the pattern clear:
@@
Participants were most accurate when no AI was involved at all (noAI/noAI), and least accurate when the idea came from AI but participants elaborated themselves (withAI/noAI). Notably, keeping the workflow consistent (withAI/withAI) substantially improved accuracy relative to the two mixed workflows (79.3% vs. 37.7% and 64.0%), indicating that mismatched sources between ideation and elaboration particularly impair memory for where the idea originated.
@@ 4.1.4. Summary
For remembering the idea source, accuracy was substantially lower. Participants were most accurate when no AI was involved. Using AI consistently was less accurate (withAI/withAI) but higher than mixed workflows: Using AI for only the idea or only the elaboration reduced the odds, with withAI/noAI being the least accurate.
For remembering the elaboration source, the ranking of conditions was identical to idea-source memory, but overall accuracy was higher.
@@ 4.2. Source Attribution Confidence
As an overview, Figure 6 shows interaction plots of model-based mean confidence in idea and elaboration attribution.
@@ 4.2.1. Confidence in Remembering the Idea Source
For confidence in remembering the idea source, the grand mean across conditions was 81.10 (95% CI [79.29, 82.91]). Confidence was highest when both idea and elaboration originated without AI support (EMM = 85.2). Confidence dropped when AI was involved in either stage (EMMs = 80.0 and 80.3) and was lowest when AI was used in both stages (78.9). A significant Idea $\times$ Elaboration interaction ($\beta=0.95$, $p<.001$) indicated that the reduction with AI involvement was not purely additive. Figure 6 (left) shows the interaction.
The model-based estimated marginal means (Table 4, from emmeans, 0–100 scale) were:
@@ 4.2.2. Confidence in Remembering the Elaboration Source
For elaboration source confidence, the grand mean across conditions was 84.0 (95% CI [82.2, 85.9]). Confidence was highest when both idea and elaboration were created without AI (EMM = 86.3). It dropped when AI was involved at either stage: ideas ($\beta=-0.82$, $p<.001$) and elaborations ($\beta=-0.63$, $p<.001$). A significant interaction ($\beta=0.81$, $p<.001$) indicated that when both stages were AI, confidence was higher than expected from adding the two main effects. Figure 6 (right) shows the interaction.
@@
Pairwise contrasts (Holm-adjusted) confirmed that the all-human workflow (noAI/noAI) elicited significantly greater confidence than any AI-involved workflow ($p<.001$). By contrast, we found not significant differences between the AI-involved workflows.
@@ 4.2.3. Metacognitive Sensitivity
Metacognitive sensitivity describes the ability to accurately evaluate one’s own decisions, thoughts, or perceptions by distinguishing between correct and incorrect responses. We operationalized this as the difference in confidence between (ground truth) correct and incorrect attributions (Correct - Incorrect), estimated with emmeans from the integrated models that included ground truth and correctness factors. Positive values indicate higher confidence when the attribution was correct.
@@
Pairwise comparisons (Holm-adjusted) showed that noAI/noAI exhibited a larger sensitivity than withAI/noAI (difference in $\Delta=9.74$, $p<.0001$), noAI/withAI ($8.92$, $p=.0001$), and withAI/withAI ($5.53$, $p=.0136$); withAI/withAI exceeded withAI/noAI ($4.21$, $p=.0366$); other differences were not significant. Thus, participants’ confidence tracked correctness most strongly without AI, moderately when using AI for both idea and elaboration, and weakly in mixed workflows. Figure 7 shows the confidence for Correct vs. Incorrect attributions within each ground truth combination.
Figure 7. Metacognitive sensitivity plots for idea: confidence for Correct vs. Incorrect attributions within each ground truth combination. Points are estimated marginal means with the same (Holm-adjusted) 95% CIs as used elsewhere; green = Correct, red = Incorrect. idea-source confidence shows a large correctness gap in noAI/noAI, a moderate gap in withAI/withAI, and minimal gaps in mixed workflows.
[…]
@@
Contrasts (Holm) indicated a larger sensitivity in withAI/withAI than noAI/noAI (difference in $\Delta=5.41$, $p=.0112$); the remaining differences among AI-involved workflows were not significant (all $p>.17$). Thus, participants’ confidence in elaboration attribution tracked correctness in all workflows, with the highest sensitivity when AI was used for both idea and elaboration. Figure 8 shows the confidence for Correct vs. Incorrect attributions within each ground truth combination.
Figure 8. Metacognitive sensitivity plots for elaboration: confidence for Correct vs. Incorrect attributions within each ground truth combination. Points are estimated marginal means with the same (Holm-adjusted) 95% CIs as used elsewhere; green = Correct, red = Incorrect. elaboration-source confidence shows positive correctness gaps in all cells, largest in withAI/withAI.
[…]
@@ 4.2.4. Confidence-Accuracy Correlation
We quantified metacognitive calibration by correlating trial-level confidence with attribution accuracy (Correct =1=1, Incorrect =0=0) using Pearson point-biserial correlations. We report overall effects and effects within each ground truth combination of idea and elaboration sources (all $p$-values after Holm correction).
@@
For idea judgments, calibration depended on the workflow: Confidence tracked accuracy in consistent workflows (noAI/noAI $r=.228$; withAI/withAI $r=.234$; both $p<.001)$, but not in mixed workflows: noAI/withAI was essentially null (r=.001r=-.001, $p=.976$), and withAI/noAI showed a small negative association (r=.094r=-.094, $p=.0011$).
@@
Table 8. Point-biserial correlations between confidence and accuracy within each GT combination. Holm-adjusted $p$-values control the familywise error rate within each outcome.
[…]
@@ 4.2.5. Summary
For this analysis, we first computed per-participant accuracy rates for both idea source and elaboration source attribution as follows: For items seen in Phase 1, a response was correct if the item was recognised as seen and the reported source matched the ground truth. For distractor items, a response was correct if participants correctly detected that they had not worked on this in Phase 1.
@@
Figure 9 plots objective accuracy against participants’ self-reports for each measure.
@@ 4.4. Multinomial Processing Tree (MPT) Model
Recognition experiments with categorical responses are better analysed with MPT models (Bröder and Meiser, 2007). For a tutorial, see Schmidt et al. (2023). An MPT specifies a small set of latent cognitive states (e.g. “detect old”, “remember source”, “guess”) and the probabilistic paths that link those states to the observed response frequencies. MPTs therefore serve as computational models of recognition memory. Compared with raw proportion tests or LMMs, MPTs disentangle cognitive memory processes from pure response and yield parameter estimates that are directly interpretable in psychological terms.
Because joint trees of considering idea and elaboration recognition together failed to reproduce the strong within-item covariance, we modelled idea and elaboration responses with two independent 2-high-threshold (2-HT) MPTs (Snodgrass and Corwin, 1988). The component-wise 2-high-threshold model assigns a single probability to each latent processing step that links the unobserved memory state to the observed categorical response. All parameters are constrained to the unit interval and are estimated hierarchically (i.e., a group mean and an across-person standard deviation). For every component, the four free parameters have the following psychological interpretation, as shown in Figure 10.
All parameters were estimated hierarchically with TreeBUGS (Heck et al., 2018) using 40 000 MCMC iterations (after 5 000 adaptation), three parallel chains, and a thinning interval of 10. TreeBUGS fits MPT in a Bayesian hierarchical framework, allowing the researcher to account for repeated measures and multiple subjects, sampling the joint posterior of individual and group parameters with JAGS and returning full uncertainty intervals plus posterior-predictive $p$-values for mean and covariance structure. We used standard priors to facilitate convergence. Gelman–Rubin diagnostics were excellent ($\hat{R}\leq 1.01$). Posterior-predictive checks showed small mean discrepancies (T1 obs–pred <5<5), but covariance misfit (T2, $p=0$) indicates residual trial-level dependence that separate trees cannot capture. Parameter estimates are nevertheless robustly addressed for our preregistered H1.

### v2

@@ 3.4.1. Phase 1 – Creating Ideas and Elaborations
The task description displayed the problem text and instructed participants to create five ideas each consisting of 1-3 keywords. These ideas were entered into five input fields displayed on one page (Fig. 3 A).
In the withAI condition, a chatbot interface was available on the right side of the screen (Fig. 3 C). Participants could copy text from the chatbot and paste it into the input fields using the usual keyboard shortcuts (Ctrl/Cmd + C, Ctrl/Cmd + V), right-click or buttons located next to each field (Fig. 3 A, on the right). It was only possible to copy text directly from our chatbot, and manual typing was restricted to deletions (i.e. with backspace) to ensure that participants directly used the AI output. In the noAI condition, the chatbot was locked (Fig. 3 D) and pasting any text into the input fields was disabled.
@@
Following this, participants were asked to elaborate on each of their previously submitted ideas with a one-sentence elaboration. Each such elaboration task displayed the task instruction, the original problem statement, and one of the participant’s earlier ideas alongside a text input field (Fig. 3 B).
@@
For both ideation and elaboration, the overall progress was shown at the bottom of the application (Fig. 3 E). After completing elaborations for all 40 ideas, the study concluded with a thank-you screen. Participants were reminded that they would receive an invitation to Phase 2 six days later and that this follow-up phase would then need to be completed within 48 hours.
@@ 3.4.2. Phase 2 - Memory and Source Attribution
After the six-day delay, participants who completed Phase 1 received an invitation to take part in Phase 2. They were presented with the study information and again provided informed consent before proceeding.
We first asked participants to guess the purpose of the study in a free text entry field. The main part of Phase 2 then consisted of the memory and source attribution tasks (Fig. 4). Besides the overall task description, each task page displayed the original problem statement and a solution (elaboration) text created by the participant in Phase 1. The underlying ideas were not shown.
Participants first indicated whether they remembered working on the displayed solution, and if they did: (1) Whether they came up with the underlying idea on their own or with AI support, (2) How confident they were in that attribution, (3) Whether they wrote the solution text on their own or with AI support, and (4) How confident they were in that second attribution. If they did not remember working on the displayed solution at all, they skipped the attribution questions and proceeded to the next memory task.
@@ 3.5. Data sharing
We make our experimental software, study material, analysis code, and the collected data accessible through our project repository: https://osf.io/dvkj9/. Two researchers went through all the data collected via free text input fields to ensure that all potentially privacy-sensitive information were removed before publication.
@@ 4.1. Source Attribution Accuracy
Fig. 5 shows the model-based interaction plots of predicted source attribution accuracy for all three memory types across conditions (noAI vs. withAI).
@@ 4.1.1. Remembering the Item
Pairwise contrasts (Holm-adjusted) showed that the no-AI workflow (noAI/noAI) yielded higher accuracy than any condition with AI: vs. withAI/noAI, OR =2.02=2.02; vs. noAI/withAI, OR =3.29=3.29; vs. withAI/withAI, OR =8.15=8.15, all $p<.001$. Among those, withAI/noAI exceeded noAI/withAI (OR =1.99=1.99) and withAI/withAI (OR =4.01=4.01); noAI/withAI also exceeded withAI/withAI (OR =2.45=2.45); again all $p<.001$. These results indicate that both AI idea generation and AI elaboration independently impair memory, and their joint use compounds this effect, with the strongest decrement observed when AI contributed at the elaboration stage.
@@ 4.1.2. Remembering the Source of Ideas
Both main effects were negative and significant, qualified by a strong positive interaction: $\beta_{\text{idea({withAI}{})}}=-3.05$, $\beta_{\text{elab({withAI}{})}}=-1.97$, $\beta_{\text{idea({withAI}{})}\times\text{elab({withAI}{})}}=3.81$; all $p<.001$. That is, using AI for the idea alone sharply reduced correct attribution (OR =0.05=0.05), as did using AI only for elaboration (OR =0.14=0.14); when both idea and elaboration came from AI, the interaction increased the odds ($\times 45.3$), partially offsetting the main-effect penalties. Model-based predicted probabilities (Table 3) make the pattern clear:
@@
Participants were most accurate without AI (noAI/noAI), and least accurate when the idea came from AI but participants elaborated themselves (withAI/noAI). Notably, keeping the workflow consistent (withAI/withAI) substantially improved accuracy relative to the two mixed workflows (79.3% vs. 37.7% and 64.0%), indicating that mismatched sources between ideation and elaboration particularly impair memory for where the idea originated.
@@ 4.1.4. Summary
For remembering the idea source, accuracy was substantially lower. Participants were most accurate when no AI was involved. Using AI consistently (withAI/withAI) was less accurate but higher than mixed workflows: Using AI for only the idea or only the elaboration reduced the odds, with withAI/noAI being the least accurate.
For remembering the elaboration source, the ranking of conditions was identical to idea source memory, but overall accuracy was higher.
@@ 4.2. Source Attribution Confidence
As an overview, Fig. 6 shows interaction plots of model-based mean confidence in idea and elaboration attribution.
@@ 4.2.1. Confidence in Remembering the Idea Source
After evaluating accuracy, we next assess participants’ confidence in idea-source attributions to see how well participants are able to judge their own ability to remember whether an idea originated from AI vs. themselves. For confidence in remembering the idea source, the grand mean across conditions was 81.10 (95% CI [79.29, 82.91]). Confidence was highest when both idea and elaboration originated without AI support (EMM = 85.2). Confidence dropped when AI was involved in either stage (EMMs = 80.0 and 80.3) and was lowest when AI was used in both stages (78.9). A significant Idea $\times$ Elaboration interaction ($\beta=0.95$, $p<.001$) indicated that the reduction with AI involvement was not purely additive. Fig. 6 (left) shows the interaction.
The estimated marginal means (Table 5, from emmeans, 0–100 scale) were:
@@ 4.2.2. Confidence in Remembering the Elaboration Source
We now analyse confidence in elaboration-source judgments to determine whether metacognitive insight differs between ideation and elaboration. For elaboration source confidence, the grand mean across conditions was 84.0 (95% CI [82.2, 85.9]). Confidence was highest when both idea and elaboration were created without AI (EMM = 86.3). It dropped when AI was involved at either stage: ideas ($\beta=-0.82$, $p<.001$) and elaborations ($\beta=-0.63$, $p<.001$). A significant interaction ($\beta=0.81$, $p<.001$) indicated that when both stages were AI, confidence was higher than expected from adding the two main effects. Fig. 6 (right) shows the interaction.
@@
Pairwise contrasts (Holm-adjusted) confirmed that the all-human workflow (noAI/noAI) elicited significantly greater confidence than any AI-involved workflow ($p<.001$). By contrast, we found no significant differences between the AI-involved workflows.
@@ 4.2.3. Metacognitive Sensitivity
Metacognitive sensitivity describes the ability to accurately evaluate one’s own decisions, thoughts, or perceptions by distinguishing between correct and incorrect responses. Here, we compare confidence for correct versus incorrect responses, allowing us to assess how well participants can monitor their memory performance in human-AI workflows. We operationalized this as the difference in confidence between (ground truth) correct and incorrect attributions (Correct - Incorrect), estimated with emmeans from the integrated models that included ground truth and correctness factors. Positive values indicate higher confidence when the attribution was correct.
@@
Pairwise comparisons (Holm-adjusted) showed that noAI/noAI exhibited a larger sensitivity than withAI/noAI (difference in $\Delta=9.74$, $p<.0001$), noAI/withAI ($8.92$, $p=.0001$), and withAI/withAI ($5.53$, $p=.0136$); withAI/withAI exceeded withAI/noAI ($4.21$, $p=.0366$); other differences were not significant. Thus, participants’ confidence tracked correctness most strongly without AI, moderately when using AI for both idea and elaboration, and weakly in mixed workflows. Fig. 7 shows the confidence for Correct vs. Incorrect attributions within each ground truth combination.
Figure 7. Metacognitive sensitivity plots for idea: confidence for Correct vs. Incorrect attributions within each ground truth combination. Points are estimated marginal means with the same (Holm-adjusted) 95% CIs as used elsewhere. Idea-source confidence shows a large correctness gap in noAI/noAI, a moderate gap in withAI/withAI, and minimal gaps in mixed workflows.
@@
Contrasts (Holm) indicated a larger sensitivity in withAI/withAI than noAI/noAI (difference in $\Delta=5.41$, $p=.0112$); the remaining differences among AI-involved workflows were not significant (all $p>.17$). Thus, participants’ confidence in elaboration attribution tracked correctness in all workflows, with the highest sensitivity when AI was used for both idea and elaboration. Fig. 8 shows the confidence for Correct vs. Incorrect attributions within each ground truth combination.
Figure 8. Metacognitive sensitivity plots for elaboration: confidence for Correct vs. Incorrect attributions within each ground truth combination. Points are estimated marginal means with the same (Holm-adjusted) 95% CIs as used elsewhere. Elaboration-source confidence shows correctness gaps in all cells, largest in withAI/withAI.
@@ 4.2.4. Confidence-Accuracy Correlation
Finally, we correlate confidence with accuracy to evaluate overall calibration and determine whether confidence reliably predicts correctness across workflows. We quantified this metacognitive calibration by correlating trial-level confidence with attribution accuracy (Correct =1=1, Incorrect =0=0) using Pearson point-biserial correlations. We report overall effects and effects within each ground truth combination of idea and elaboration sources (all $p$-values after Holm correction).
@@
For idea judgments, calibration depended on the workflow: Confidence tracked accuracy in consistent workflows (noAI/noAI $r=.228$; withAI/withAI $r=.234$; both $p<.001)$, but not in mixed workflows: noAI/withAI was essentially zero (r=.001r=-.001, $p=.976$), and withAI/noAI showed a small negative association (r=.094r=-.094, $p=.0011$).
@@
Table 9. Point-biserial correlations between confidence and accuracy in each GT combination. Holm-adjusted $p$-values control the familywise error rate within each outcome.
@@ 4.2.5. Summary
To examine how well participants evaluated their own memory performance, we compared their actual source-attribution accuracy with their self-assessed performance. This reveals whether participants’ self-evaluations aligned with their actual performance or showed systematic misjudgment. For this analysis, we first computed per-participant accuracy rates for both idea source and elaboration source attribution as follows: For items seen in Phase 1, a response was correct if the item was recognised as seen and the reported source matched the ground truth. For distractor items, a response was correct if participants correctly detected that they had not worked on this in Phase 1.
@@
Fig. 9 plots objective accuracy against participants’ self-reports for each measure.
@@ 4.4. Multinomial Processing Tree (MPT) Model
Recognition experiments with categorical responses are better analysed with MPT models (Bröder and Meiser, 2007). For a tutorial, see Schmidt et al. (2023). An MPT specifies a small set of latent cognitive states (e.g. “detect old”, “remember source”, “guess”) and the probabilistic paths that link those states to the observed response frequencies. We use “old” and “new” in a standard recognition-memory sense: old items are the solution texts that participants had previously created in Phase 1, whereas new items refer to the distractors (known-topic or unknown-topic), only shown in Phase 2. MPTs thus serve as computational models of recognition memory. Compared with raw proportion tests or LMMs, MPTs disentangle cognitive memory processes from pure response and yield parameter estimates that are directly interpretable in psychological terms. In line with standard applications of the MPT, our model assumes that, within each component (idea vs elaboration), each participant is characterised by a set of latent probabilities ($d,s,\beta,f$) that apply across the set of items in that component. Thus, the MPT does not attempt to model a single solution separately but rather captures each participant’s average detection, source-memory, and guessing tendencies over the specific ideas and elaborations included in the task.
Because joint trees of considering idea and elaboration recognition together failed to reproduce the strong within-item covariance, we modelled idea and elaboration responses with two independent 2-high-threshold (2-HT) MPTs (Snodgrass and Corwin, 1988). The component-wise 2-high-threshold model assigns a single probability to each latent processing step that links the unobserved memory state to the observed categorical response. Practically, this means that for a given participant and component (idea vs elaboration), the model assumes a probability of detecting seen items ($d$), a probability of remembering its source conditional on detection ($s$), a probability of guessing “with AI” when the source is not remembered ($\beta$) and one false alarm rate for distractors ($f$), averaged across all items of that type. All parameters are constrained to the unit interval and are estimated hierarchically (i.e., a group mean and an across-person standard deviation). For every component, the four free parameters have the psychological interpretation shown in Fig. 10.
All parameters were estimated hierarchically with TreeBUGS (Heck et al., 2018) using 40 000 MCMC iterations (after 5 000 adaptation), three parallel chains, and a thinning interval of 10. TreeBUGS fits MPT in a Bayesian hierarchical framework, allowing the researcher to account for repeated measures and multiple subjects, sampling the joint posterior of individual and group parameters with JAGS and returning full uncertainty intervals plus posterior-predictive $p$-values for mean and covariance structure. We used standard priors to facilitate convergence. Gelman–Rubin diagnostics were excellent ($\hat{R}\leq 1.01$). Posterior-predictive checks showed small mean discrepancies (T1 obs–pred <5<5), but covariance misfit (T2, $p=0$) indicates residual trial-level dependence that separate trees cannot capture. This aligns with the idea that individual solutions vary in how diagnostic their style and content are for identifying the source (see Section 4.5.2). Consequently, the MPT parameters should be interpreted as summaries of average behaviour across the item set rather than precise probabilities for each individual solution. Parameter estimates are nevertheless robustly addressed for our preregistered H1.